# Threshold model of language competition including the bilingual state

### v1

@@ Abstract
We propose a threshold model of language competition which includes intermediate bilingual state. The model is based on the Minett-Wang model , but through the introduction of thresholds in the language shift rates, it incorporates the effects of memory and learning. The model is piecewise-linear, allowing the exact analytical treatment. We study the symmetric case where two competing languages are equivalent in terms of status and social pressure and provide a complete list of the various dynamical regimes. We also study several limiting regimes corresponding to asymmetric systems and characterize the full spectrum of possible asymptotic behaviors. Unlike the Minett-Wang model, which always predicts the extinction of one of the languages, the proposed new model exhibits a wide range of possible equilibrium scenarios, including equilibrium states of coexistence. Most commonly, in such coexistence regimes the minority language speakers are either completely monolingual or completely bilingual.
@@ II Three-state models with bilinguals
Figure 1: Panels (A) and (B) represent the state of a three-state system in Cartesian coordinates and by a tertiary diagram, respectively. Panel (C): learning and attrition thresholds (4b),(5b) correspond to straight lines, each of them splits the ternary diagram into regions where the corresponding process is on or off; Panel (D): whenever parameters $\beta^{+}$ and $\beta^{-}$ are unequal, the learning and retention thresholds may intersect; if $\beta^{-}>\beta^{+}$ the steeper line corresponds to the retention threshold.
@@ II.2 The role of bilinguals in learning and attrition
MW model described above implicitly makes a particular choice concerning the role of bilinguals in language learning and attrition. To formulate a more general three-state model it is instructive to discuss this role here explicitly.
@@ II.3 Threshold model
Figure 2: Comparison of the X-language normalized rates of (A) learning, $r_{\scriptscriptstyle Y\to Z}/k_{YZ}$, and (B) attrition, $r_{\scriptscriptstyle Z\to Y}/k_{ZY}$, versus the corresponding effective fraction of X-speakers of the language to be learned, $x_{\mathrm{eff}}^{+}$, or maintained, $x_{\mathrm{eff}}^{-}$, respectively, for different three-state models (a similar scheme holds for the Y-language): MW bilinguals model with volatility $a>1$, Eqs. (5) (dashed black curves); threshold model with step function rates, Eqs. (6) (thin blue solid curves); and heterogeneous threshold model with sigmoidal shape (thick red solid curves).
@@
where $\Theta(x)$ is the Heaviside theta-function, the effective fractions of $x,y$ speakers $x_{\mathrm{eff}}^{\pm},y_{\mathrm{eff}}^{\pm}$ are given by (4b),(5b) and the parameters $x_{\pm}^{*},\,y_{\pm}^{*}$ represent their threshold values. Each of the conditions at which Heaviside functions in (6) switch from zero to one,
@@
Figure 3: Dynamics of system (10) in the case when each reaction channel is either permanently open (red arrow) or permanently closed (gray arrow). The bold dots and lines on the ternary plots indicate the attractors, thin line a schematic representation of evolution trajectories, arrows on them indicate the direction of the evolution, gray area is the area of neutral stability. The exemplary trajectories are drawn for a fully symmetric case $\gamma_{x}=\gamma_{y},\kappa=1/2$. The last column introduces color-coding of the states, which is used throughout the following pictures. All states except 1, 3, 7 and 10 have dual counterparts, which can be obtained by $X\longleftrightarrow Y$ replacement and are not shown for brevity.
@@ III.1 Pure regimes
At any time $t$, each Heaviside function in Eqs. (10), depending on the values of the population fractions $x(t),y(t),z(t)$ can take either value 0 or 1, meaning that the correspondent process is switch off or on, respectively. Altogether, for each combination of $x,y,z$ there are $2^{4}=16$ combinations of possible values of the Heaviside functions and therefore 16 different possible dynamical regimes. Generally, a trajectory of the system is associated with more than one of these dynamical regimes: while $x(t)$ and $y(t)$ change, the values of corresponding Heaviside functions might switch between 0 and 1.
However, it is instructive to start the study of Eqs. (10) by considering the situations where each of the 4 Heaviside functions is resolved in the same way for all values of $x,y,z$, i.e., where all four functions are always 0, where the first one is always 1 and the other three are always 0, etc. We refer to such situations as “pure regimes” of Eqs. (10), as opposed to the more general mixed regimes, where Heaviside functions are resolved differently for different values of $x,y,z$. Indeed, the pure dynamic regimes essentially are the building blocks from which the solutions of mixed regimes can be constructed.
@@ III.1.4 Intermediate regimes
Figure 4: Dynamics of an exemplary system with $x^{*}_{+}=y^{*}_{+}=0.3,x^{*}_{-}=y^{*}_{-}=0.15,\kappa=1/2$. (A) the switching conditions split the whole space of possible compositions into 7 areas, numbers and colors correspond to the pure states as introduced in Fig. 3 and in the text; (B) schematic depiction of the dynamic flows in the system, red dashed lines separate basins of attraction; (C) attractors (black bold lines and points) of Eqs. (10) for this particular set of parameters and boundaries of their basins of attraction; red dots and lines correspond to the initial conditions and trajectories of the dynamics shown in panels (D)-(F); (D)-(F) exemplary dynamics for 3 different sets of initial conditions for $\gamma_{x,y}=0.1$; red, green and orange lines show the time evolutions of $x(t),y(t),z(t)$ given by (10) with Heaviside (full lines) and S-shaped ((34), dashed lines); the dashed vertical lines indicate the switching between the regimes, with the numbers inside the panels indicating the regimes; (D) initial condition $(x_{0},y_{0},z_{0})=(0.75,0.25,0)$ converging to the full extinction of language Y, $(x_{\infty},y_{\infty},z_{\infty})=(1,0,0)$; (E) evolution starting from initial condition $(x_{0},y_{0},z_{0})=(0.6,0.4,0)$ converging to to a long-term coexistence of monolinguals of X and bilinguals , $(x_{\infty},y_{\infty},z_{\infty})=(0.45,0.55,0)$;, (F) evolution starting from initial condition $(x_{0},y_{0},z_{0})=(0.1,0.15,0.75)$ is frozen from the start, so $(x_{\infty},y_{\infty},z_{\infty})=(0.1,0.15,0.75)$.
@@ III.2 An example of the dynamics
Figure 5: Flows and attractors of the system with no language attrition. Top row $x^{*}_{+}+y^{*}_{+}>1$, bottom row $x^{*}_{+}+y^{*}_{+}\leq 1$. First column shows splitting of the space of parameters into regions where the conditions in Eqs. (35) are opened differently, second column is a flaw diagram, third column shows the attractors of the dynamics. Note that the shape of the vertical line splitting the basins of attraction in the bottom raw is, generally speaking, $\kappa$-dependent and is a straight line going through $z=1$ angle of the ternary diagram only if $\kappa=1/2$. The gray point inside the triangle in the bottom row corresponds to a partially stable stationary point at $(x,y,z)=(x_{+}^{*},y_{+}^{*},1-x^{*}_{+}-y_{+}^{*})$, it is attractive along the vertical black line, repulsive in perpendicular direction, neutral if approached from the gray area side.
As an example, consider a fully symmetric system with $\kappa=1/2$ $x^{*}_{+}=y^{*}_{+}=0.3$, $x^{*}_{-}=y^{*}_{-}=0.15$. This means that if more than 30% of the population are monolinguals of one or the other language, this language will be learned by the monolinguals of the other language, but if the effective fraction of the speakers of one of the languages is below 15%, this language will be forgotten by bilinguals [see Eqs. (LABEL:eq:xy_eff), (6), (8)]. The conditions at which the learning and forgetting processes switch on,
@@
In Fig. 4 D-F with dashed lines we show the numerical solutions of (10) with such S-shaped rates for $w=0.1$. In the extinction regime (Fig. 4D) replacing the theta-functional rates with S-shaped ones leads only to minor quantitative changes in the dynamic. In the cases when theta-functional rates lead to fully (Fig. 4F) or partly (Fig. 4E) frozen dynamics the changes are more profound. In both cases, the dynamics with S-shaped rates leads to the eventual extinction of the minority language, but the time it takes is of order hundreds, i.e., well above the observational timescales (recall that the unite of time is around 10 years). Meanwhile, stable states of the system with theta-functional thresholds survive as long-living intermediate states in the case of S-shaped ones.
Indeed, in Fig. 4E the coexistence between monolinguals of the majority language and bilinguals, which survives ad infinitum for theta-functional rates, exists as an intermediate states for $5\lesssim t\lesssim 50$, i.e., for several centuries. In Fig. 4F the evolution is essentially frozen up to $t\approx 30$, then monolinguals of the minority language (X in this case) disappear and for $30\lesssim t\lesssim 300$ there is a long-living intermediate state consisting of bilinguals and monolinguals Y.
@@ III.3 Limiting cases
Figure 6: Flows and attractors of the system where language attrition is always on. Three rows correspond to three distinct cases with respect to the values of $x^{*}_{+},y^{*}_{+}$. First row: $x^{*}_{+}+y^{*}_{+}>1$. Second and third rows: $x^{*}_{+}+y^{*}_{+}\leq 1$. Third row corresponds to the case when stable point is inside the dark purple triangle, i.e. $x^{*}_{+}>x_{\infty},y^{*}_{+}>y_{\infty}$, see (37), second raw – to the case when it is not, i.e., either $x^{*}_{+}<x_{\infty}$ or $y^{*}_{+}<y_{\infty}$. Similarly to Fig. 5 the columns correspond to (i) splitting of the space of parameters into regions, (ii) flaw diagram, and (iii) attractors and basins of attraction. The slope of trajectories in the red region 3 is determined by the ratio $\gamma_{x}/\gamma_{y}$ (vertical if it equals 1). In a fully symmetric region the line separating basins of attraction in the second row is vertical. The gray dot in this diagram designates a saddle point at $(x,y,z)=(x_{+}^{*},y_{+}^{*},1-x^{*}_{+}-y_{+}^{*})$.
@@
Figure 8: The plane of parameters $0<s_{+}^{*}<1$ $0<s_{-}^{*}<1$ split into subregions with different topology of theta-function conditions intersections, corresponding to different dynamic regimes. See Fig. 9 for ternary diagrams corresponding to each subregion.
Figure 9: Topology of the conditions overlap and corresponding steady states and basins of attraction for each region in diagram in Fig. 8. First column: region on the $(x_{+}^{*},x_{-}^{*})$ plane, according to numbering in Fig. 8. Second column: the ternary X-Y-Z diagram with the position of the lines defined by the conditions in (11). The areas are color-coded according to corresponding pure states (see Fig. 3). Third column: attractors of the dynamics and their corresponding basins of attraction: black points correspond to attractive stable points, thick lines - to stable points neutral in one direction (along the line) and attractive in the perpendicular one; gray areas - to stable points neutral in both direction. Saddle points and other partially repulsive stable points are not shown are not shown in order not to overcomplicate the picture. Thin line schematically delineate the borders of basins of attraction.
Second, if $x^{*}_{+}+y^{*}_{+}>1$ the outcome additionally depends on whether the attractive point of regime 10, corresponding to Eqs. (23), is inside or outside the corresponding area (dark purple triangle) in Fig. 6. If it is outside, as shown in the second row of Fig. 6, then all trajectories starting in this region eventually end up on its boundary, switching to regime 7. As a result, in this case the only possible long-term outcome is the extinction of one of the languages (except for a saddle point at $x=x_{+}^{*}$, $y=y_{+}^{*}$, which is attractive when approached from regions of regimes 3 and 10 and repulsive when approached from regions or regime 7). Conversely, if the stable point is inside the triangular region, a stable multilingual equilibrium becomes possible. The condition for this to happen reads
@@ III.4 Fully symmetric case
I. Full extinction of one of the languages; these are two attractive stable points at $(x,y,z)=(1,0,0)$ and $(x,y,z)=(0,1,0)$, marked by heavy dots at the X and Y corners of the tertiary diagrams.
@@ IV Discussion and Conclusion
In this paper we generalized the threshold model of language competition introduced in Ref. Tamm-2025 to the three-state case where the existence of bilinguals is taken into account explicitly. The model includes learning and attrition thresholds that encode the minimal exposure needed to acquire or maintain a language, and allows bilinguals to influence learning and attrition asymmetrically via different effective fractions of speakers for these two processes. The suggested threshold-based formulation ensures that the dynamics of language competition is described by a system of piecewise-linear differential equations, which are analytically tractable. Solution of the model consists of two stages. First, we solve the system for each possible combination of values of $\Theta$-function in Eqs. (10), getting the solutions of what we call “pure regimes”. At any given instance of time the actual solution of Eqs. (10) belongs to one of these regimes. Second, these solutions need to be glued together at points where variables $x$ and $y$ cross a threshold and the corresponding $\Theta$-function switches either on or off. As a result for any given set of parameters the full solution is a concatenation of the pure regime solutions.
@@
Moreover, there are two outcomes similar to those we predicted in the analogous two-state model Tamm-2025: the completely frozen state, and the attractive stable point corresponding to stable language coexistence.
@@
Fourth, and most interestingly, there is a new class of mixed outcomes, corresponding to partially attractive and partially neutral stable points, i.e., stable points which form one-dimensional manifolds so that, while the outcomes are stable with respect to perturbation of the initial condition in one direction, perturbation in the other directions results in shifting the outcome along this manifold of stable points. In the symmetric or weakly asymmetric systems there are two examples of such mixed outcomes. The first one (regime IV of the symmetric system) corresponds to a split of the society into two non-interacting monolingual communities. This state is attractive in the sense that bilinguals initially present in the system become extinct, and neutral in that the exact proportion in which the population is split between the two languages depends on the initial conditions. The second possibility (regime III of the symmetric system) is an equilibrium in which the monolingual sub-population of one language (say, X) remains monolingual, while all monolinguals of the language Y become bilingual. In this case, the equilibrium is stable with respect to changing the initial composition between monolinguals Y and bilinguals, but the exact split between monolingual X and bilingual communities in the final state depends on initial conditions.
@@
This result is obtained under the assumption that there are ideally sharp theta-functional thresholds in the learning and attrition rates. If actual rates are more smeared S-shaped function of effective densities (see red lines in Fig. 2) we expect, similarly to the two-state model, that the frozen state will degrade eventually but at a time scale that far exceeds the typical observational scales. In turn, intermediate states turn into long-living transient ones: first the system converges at a time scale of order at most $\gamma^{-1}$, then these states degrade towards attractive stable points (language extinction or coexistence) at much larger time scale controlled by a particular form of the S-shaped rates.
@@
Such long-lived intermediate states should be observed in reality: bilingual societies should disproportionally often be observed in the vicinity of one of the mixed stable points. That is to say, minority language communities should often be either predominantly functionally monolingual and thus separated from the majority (i.e., close to intermediate regime IV) or predominantly functionally bilingual (i.e., close to intermediate regime III), while situations where the minority language speakers are split into comparable number of bilinguals and monolinguals should be relatively rare. As we show in Table 1, which uses the data from Sanchez-2021, this indeed seems to be the case: out of 12 societies with two dominant languages, 3/4 demonstrate either clear domination of monolingualism among minority-language speakers (e.g., in Belgium or Estonia), or clear domination of bilingualism (e.g., in Galicia and Basque Country). Note that Twitter data is, clearly, a skewed sample of the population (e.g., probability of using Twitter might be language-dependent). However, we expect it to provide a good qualitative insight into dominance of monolingual and bilingual behavior within the minority community. Although a detailed comparison with the historical dynamics of language shift in these societies is impossible due to absence of long-term quality data, we believe that this qualitative observation is an important argument in support of the presence of intermediate semi-stable states predicted in this paper.
@@
In the future, it is interesting to study how thresholds in language learning and attrition might interplay with spatial inhomogeneities, demographic processes and network structure of society to give rise to further complexity in the dynamics of language shift, supporting data-driven calibration and policy counterfactuals (e.g., how changes in exposure or retention environments reshape the landscape of outcomes). Beyond language dynamics, the same thresholded three-state formalism should be applicable to other social and possibly biological systems where acquisition and maintenance are driven by exposure-dependent processes.

### v2

@@ Abstract
We propose a threshold model of language competition which includes an intermediate bilingual state. The model is based on the Minett-Wang model but through the introduction of thresholds in the language shift rates, it incorporates the effects of memory and learning. The model is piecewise-linear, allowing an exact analytical treatment. We study the symmetric case where two competing languages are equivalent in terms of status and social pressure and provide a complete list of the various dynamical regimes. We also study several limiting regimes corresponding to asymmetric systems and characterize the full spectrum of possible asymptotic behaviors. Unlike the Minett-Wang model, which always predicts the extinction of one of the languages, the proposed new model exhibits a wide range of possible equilibrium scenarios, including equilibrium states of coexistence. Most commonly, in such coexistence regimes the minority language speakers are either predominantly monolingual or predominantly bilingual.
@@ II Three-state models with bilinguals
Figure 1: Panels (A) and (B) represent the state of a three-state system in Cartesian coordinates and by a ternary diagram, respectively. Panel (C): learning and attrition thresholds (4b) and (5b) correspond to straight lines, each of them splits the ternary diagram into regions where the corresponding process is on or off; Panel (D): whenever parameters $\beta^{+}$ and $\beta^{-}$ are unequal, the learning and retention thresholds may intersect; if $\beta^{-}>\beta^{+}$ the steeper line corresponds to the retention threshold.
@@ II.2 The role of bilinguals in learning and attrition
The MW model described above implicitly makes a particular choice concerning the role of bilinguals in language learning and attrition. To formulate a more general three-state model it is instructive to discuss this role here explicitly.
@@ II.3 Threshold model
Figure 2: Panel (A): normalized learning rate, $r_{\scriptscriptstyle Y\to Z}/k_{YZ}$, versus the corresponding effective fraction of X-speakers, $x_{\mathrm{eff}}^{+}$, needed to learn language X. Panel (B): normalized attrition rate, $r_{\scriptscriptstyle Z\to Y}/k_{ZY}$, versus the corresponding effective fraction of X-speakers, $x_{\mathrm{eff}}^{-}$, needed to maintain language X. Different curves correspond to different three-state models: dashed black curves—MW bilinguals model with volatility $a>1$, Eqs. (5); thin blue solid curves—threshold model with step function rates, Eqs. (6); thick red solid curves—heterogeneous threshold model with sigmoidal shape. A similar scheme holds for the Y-language.
@@
where $\Theta(x)$ is the Heaviside $\Theta$-function, the effective fractions of $x,y$ speakers $x_{\mathrm{eff}}^{\pm},y_{\mathrm{eff}}^{\pm}$ are given by (4b),(5b) and the parameters $x_{\pm}^{*},\,y_{\pm}^{*}$ represent their threshold values. Each of the conditions at which Heaviside functions in (6) switch from zero to one,
@@
Figure 3: Dynamics of system (10) in the case when each reaction channel is either permanently open (red arrow) or permanently closed (gray arrow). The bold dots and lines on the ternary plots indicate the attractors. Thin lines correspond to a schematic representation of evolution trajectories and arrows on them indicate the direction of the evolution. Gray area is the area of neutral stability. The exemplary trajectories are drawn for a fully symmetric case $\gamma_{x}=\gamma_{y}$, $\kappa=1/2$. The last column introduces color-coding of the states, which is used throughout the following figures. All states except 1, 3, 7 and 10 have dual counterparts, which can be obtained by $X\longleftrightarrow Y$ replacement and are not shown to save the space.
@@ III.1 Pure regimes
At any time $t$, each Heaviside function in Eqs. (10), depending on the values of the population fractions $x(t),y(t),z(t)$ can take either value 0 or 1, meaning that the correspondent process is switched off or on, respectively. Altogether, for each combination of $x,y,z$ there are $2^{4}=16$ combinations of possible values of the Heaviside functions and therefore 16 different possible dynamical regimes. Generally, a trajectory of the system is associated with more than one of these dynamical regimes: while $x(t)$ and $y(t)$ change, the values of corresponding Heaviside functions might switch between 0 and 1.
However, it is instructive to start the study of Eqs. (10) by considering the situations in which each of the four Heaviside functions present in Eqs. (10) is resolved in the same way for all values of $x,y,z$, i.e., where all four functions are always 0, where the first one is always 1 and the other three are always 0, etc. We refer to such situations as “pure regimes” of Eqs. (10), as opposed to the more general mixed regimes, where Heaviside functions are resolved differently for different values of $x,y,z$. Indeed, the pure dynamic regimes essentially are the building blocks from which the solutions of mixed regimes can be constructed.
@@ III.1.4 Intermediate regimes
Figure 4: Example of dynamics with $x^{*}_{+}=y^{*}_{+}=0.3,x^{*}_{-}=y^{*}_{-}=0.15,\kappa=1/2$. (A) The switching conditions split the whole space of possible compositions into 7 areas – numbers and colors correspond to the pure states introduced in Fig. 3 and in the text; (B) schematic depiction of the dynamic flows in the system – red dashed lines separate basins of attraction; (C) attractors (black bold lines and points) of Eqs. (10) for this particular set of parameters and boundaries of their basins of attraction; red dots and lines correspond to the initial conditions and trajectories of the dynamics shown in panels (D)-(F); (D)-(F) examples of dynamics for three different sets of initial conditions for $\gamma_{x,y}=0.1$; red, green and orange lines show the time evolutions of $x(t),y(t),z(t)$ given by (10) with Heaviside (full lines) and S-shaped ((34), dashed lines) rates; the dashed vertical lines indicate the switching between the regimes, with the numbers inside the panels indicating the regimes; (D) initial condition $(x_{0},y_{0},z_{0})=(0.75,0.25,0)$ converging to the extinction of language Y, $(x_{\infty},y_{\infty},z_{\infty})=(1,0,0)$; (E) evolution starting from initial condition $(x_{0},y_{0},z_{0})=(0.6,0.4,0)$ converging to a long-term coexistence of monolinguals of X and bilinguals , $(x_{\infty},y_{\infty},z_{\infty})=(0.45,0.55,0)$;, (F) evolution starting from initial condition $(x_{0},y_{0},z_{0})=(0.1,0.15,0.75)$ is frozen from the start, so $(x_{\infty},y_{\infty},z_{\infty})=(0.1,0.15,0.75)$.
@@ III.2 An example of the dynamics
Figure 5: Flows and attractors of the system with no language attrition. Top row: $x^{*}_{+}+y^{*}_{+}>1$; bottom row $x^{*}_{+}+y^{*}_{+}\leq 1$. First column shows splitting of the space of parameters into regions where the conditions in Eqs. (35) are opened differently; second column is a flaw diagram; third column shows the attractors of the dynamics. Note that the shape of the line splitting the basins of attraction in the bottom raw is, generally speaking, $\kappa$-dependent and it is a straight line going through $z=1$ angle of the ternary diagram only if $\kappa=1/2$. The gray point inside the triangle in the bottom row corresponds to a partially stable stationary point at $(x,y,z)=(x_{+}^{*},y_{+}^{*},1-x^{*}_{+}-y_{+}^{*})$, it is attractive along the vertical black line, repulsive in perpendicular direction, neutral if approached from the gray area side.
As an example, consider a fully symmetric system with $\kappa=1/2$, $x^{*}_{+}=y^{*}_{+}=0.3$, $x^{*}_{-}=y^{*}_{-}=0.15$. This means that if more than 30% of the population are monolinguals of one or the other language, this language will be learned by the monolinguals of the other language, but if the effective fraction of the speakers of one of the languages is below 15%, this language will be forgotten by bilinguals [see Eqs. (LABEL:eq:xy_eff), (6), (8)]. The conditions at which the learning and forgetting processes switch on,
@@
In Fig. 4 D-F with dashed lines we show the numerical solutions of (10) with such S-shaped rates for $w=0.1$. In the extinction regime (Fig. 4D) replacing the $\Theta$-functional rates with S-shaped ones leads only to minor quantitative changes in the dynamic. In the cases when $\Theta$-functional rates lead to fully (Fig. 4F) or partly (Fig. 4E) frozen dynamics the changes are more profound. In both cases, the dynamics with S-shaped rates leads to the eventual extinction of the minority language, but the time it takes is of order hundreds, i.e., well above the observational timescales (recall that the unite of time is around 10 years). Meanwhile, stable states of the system with $\Theta$-functional thresholds survive as long-living intermediate states in the case of S-shaped ones.
Indeed, in Fig. 4E the coexistence between monolinguals of the majority language and bilinguals, which survives ad infinitum for $\Theta$-functional rates, exists as an intermediate states for $5\lesssim t\lesssim 50$, i.e., for several centuries. In Fig. 4F the evolution is essentially frozen up to $t\approx 30$, then monolinguals of the minority language (X in this case) disappear and for $30\lesssim t\lesssim 300$ there is a long-living intermediate state consisting of bilinguals and monolinguals Y.
@@ III.3 Limiting cases
Figure 6: Flows and attractors of the system where language attrition is always on. The three rows correspond to three distinct cases with respect to the values of $x^{*}_{+},y^{*}_{+}$. First row: $x^{*}_{+}+y^{*}_{+}>1$; second and third rows: $x^{*}_{+}+y^{*}_{+}\leq 1$. The third row corresponds to the case when the stable point is inside the dark purple triangle, i.e. $x^{*}_{+}>x_{\infty},y^{*}_{+}>y_{\infty}$, see Eqs. (37), the second row – to the case when it is not, i.e., either $x^{*}_{+}<x_{\infty}$ or $y^{*}_{+}<y_{\infty}$. Similarly to Fig. 5 the columns correspond to (i) splitting of the space of parameters into regions, (ii) flaw diagram, and (iii) attractors and basins of attraction. The slope of trajectories in the red region 3 is determined by the ratio $\gamma_{x}/\gamma_{y}$ (vertical if it equals 1). In a fully symmetric region the line separating basins of attraction in the second row is vertical. The gray dot in this diagram designates a saddle point at $(x,y,z)=(x_{+}^{*},y_{+}^{*},1-x^{*}_{+}-y_{+}^{*})$.
@@
Figure 8: The plane of parameters $0<s_{+}^{*}<1$, $0<s_{-}^{*}<1$ split into subregions with different topology of $\Theta$-function conditions intersections, corresponding to different dynamic regimes. The ternary diagrams corresponding to each subregion can be seen in Fig. 9.
Figure 9: List of topologically different ways of how the conditions under the $\Theta$-functions of Eqs. (11) may overlap and corresponding steady states and basins of attraction for each region of Fig. 8. First column: region on the $(s_{+}^{*},s_{-}^{*})$ plane, according to numbering in Fig. 8. Second column: the ternary X-Y-Z diagram with the position of the lines defined by the conditions in Eqs. (11). The areas are color-coded according to corresponding pure states (see Fig. 3). Third column: attractors of the dynamics and their corresponding basins of attraction: black points represent attractive stable points; thick lines – to stable points neutral in one direction (along the line) and attractive in the perpendicular one; gray areas – to stable points neutral in both direction. Saddle points and other partially repulsive stable points are not shown in order not to overcomplicate the picture. Thin lines schematically delineate the borders of basins of attraction.
Second, if $x^{*}_{+}+y^{*}_{+}>1$ the outcome additionally depends on whether the attractive point of regime 10, corresponding to Eqs. (23), is inside or outside the corresponding area (dark purple triangle) in Fig. 6. If it is outside, as shown in the second row of Fig. 6, then all trajectories starting in this region eventually end up on its boundary, switching to regime 7. As a result, in this case the only possible long-term outcome is the extinction of one of the languages (except for a saddle point at $x=x_{+}^{*}$, $y=y_{+}^{*}$, which is attractive when approached from regions of regimes 3 and 10 and repulsive when approached from regions or regime 7). Conversely, if the stable point is inside the triangular region, a stable multilingual equilibrium becomes possible. The condition for this to happen reads,
@@ III.4 Fully symmetric case
I. Full extinction of one of the languages; these are two attractive stable points at $(x,y,z)=(1,0,0)$ and $(x,y,z)=(0,1,0)$, marked by heavy dots at the X and Y corners of the ternary diagrams.
@@ IV Discussion and Conclusion
In this paper we generalized the threshold model of language competition introduced in Ref. Tamm-2025 to the three-state model where the existence of bilinguals is taken into account explicitly. The model includes learning and attrition thresholds that encode the minimal exposure needed to acquire or maintain a language, and allows bilinguals to influence learning and attrition asymmetrically via different effective fractions of speakers for these two processes. The suggested threshold-based formulation ensures that the dynamics of language competition is described by a system of piecewise-linear differential equations, which are analytically tractable. Solution of the model consists of two stages. First, we solve the system for each possible combination of values of $\Theta$-function in Eqs. (10), getting the solutions of what we call “pure regimes”. At any given instance of time the actual solution of Eqs. (10) belongs to one of these regimes. Second, these solutions need to be glued together at points where variables $x$ and $y$ cross a threshold and the corresponding $\Theta$-function switches either on or off. As a result for any given set of parameters the full solution is a concatenation of the pure regime solutions.
@@
There are also two outcomes similar to those we predicted in the analogous two-state model Tamm-2025: the completely frozen state, and the attractive stable point corresponding to stable language coexistence.
@@
Furthermore, there is a new class of mixed outcomes, corresponding to partially attractive and partially neutral stable points, i.e., stable points which form one-dimensional manifolds so that, while the outcomes are stable with respect to perturbation of the initial condition in one direction, perturbation in the other directions results in shifting the outcome along this manifold of stable points.
In the symmetric or weakly asymmetric systems there are two examples of such mixed outcomes. The first one (regime IV of the symmetric system) corresponds to a split of the society into two non-interacting monolingual communities. This state is attractive in the sense that bilinguals initially present in the system become extinct, and neutral in that the exact proportion in which the population is split between the two languages depends on the initial conditions. The second possibility (regime III of the symmetric system) is an equilibrium in which the monolingual sub-population of one language (say, X) remains monolingual, while all monolinguals of the language Y become bilingual. In this case, the equilibrium is stable with respect to changing the initial composition between monolinguals Y and bilinguals, but the exact split between monolingual X and bilingual communities in the final state depends on initial conditions.
@@
This result is obtained under the assumption that there are ideally sharp $\Theta$-functional thresholds in the learning and attrition rates. If actual rates are more smeared S-shaped function of effective densities (see red lines in Fig. 2) we expect, similarly to the two-state model, that the frozen state will degrade eventually but at a time scale that far exceeds the typical observational scales. In turn, intermediate states turn into long-living transient ones: first the system converges at a time scale of order at most $\gamma^{-1}$, then these states degrade towards attractive stable points (language extinction or coexistence) at much larger time scale controlled by a particular form of the S-shaped rates.
@@
Such long-lived intermediate states should be observed in reality: bilingual societies should disproportionally often be observed in the vicinity of one of the mixed stable points. That is to say, minority language communities should often be either predominantly functionally monolingual and thus separated from the majority (i.e., close to intermediate regime IV) or predominantly functionally bilingual (i.e., close to intermediate regime III), while situations where the minority language speakers are split into comparable number of bilinguals and monolinguals should be relatively rare. As we show in Table 1, which uses the data from Ref. Sanchez-2021, this indeed seems to be the case: out of 12 societies with two dominant languages, 3/4 demonstrate either clear domination of monolingualism among minority-language speakers (e.g., in Belgium or Estonia), or clear domination of bilingualism (e.g., in Galicia and Basque Country). Note that Twitter data is, clearly, a skewed sample of the population (e.g., probability of using Twitter might be language-dependent). However, we expect it to provide a good qualitative insight into dominance of monolingual and bilingual behavior within the minority community. Although a detailed comparison with the historical dynamics of language shift in these societies is impossible due to absence of long-term quality data, we believe that this qualitative observation is an important argument in support of the presence of intermediate semi-stable states predicted in this paper.
@@
In the future, it would be interesting to study how thresholds in language learning and attrition might interplay with spatial inhomogeneities, demographic processes and network structure of society to give rise to further complexity in the dynamics of language shift, supporting data-driven calibration and policy counterfactuals, e.g., how changes in exposure or retention environments reshape the landscape of outcomes. Beyond language dynamics, the same thresholded three-state formalism should be applicable to other social and possibly biological systems where acquisition and maintenance are driven by exposure-dependent processes.